I saw this problem worked out :
104% = 1.04 = 104/100 = 26/25.
My question is this: Why can 1.04 be converted to 104/100? It's the second step that confuses me.

Answer by Gogonati(855)

A number doesn't change if we multiply and divide it with the same number
for example 3=3*2/2 In our case 1.04=1.04*100/100=104/100.